Reports 1-1 of 1 Clear search Modify search
MIF (Noise Budget)
Carl Blair - 17:08 Wednesday 05 February 2025 (32591) Print this report
Comment to Noise projection from IMMT1 trans QPD to DARM (preliminary) (32588)

I projected intensity and frequency noise at the time of the transfer function measurement in to DARM with transfer functions I found in the Noisebudget TFs folder /kagra/Dropbox/Measurements/NoiseBudget/TFs/2023/0621/ : CARM2DARM_202306211610.xml and ISS2DARM_202306211610.xml.  Frequency noise produced by MCO drive and coupled in to DARM is of the correct order of magnitude to explain the observed coupling to DARM.  The shape however is not correct.  Could CARM loop or CARM coupling have achanged since June 2023.

 

Images attached to this comment
Search Help
×

Warning

×